A current 5A passes through a straight wire in a uniform magnitude field of a flux density of 2.0x10-3 T. Calculate the force per unit length exerted on the wire when it is inclined at 30° to the field.
1
Expert's answer
2020-05-18T10:01:51-0400

given data :


"I = 5A"


"B = 2.0\times 10 -3 T = 20-3 =17 T = 17 \frac{Wb}{m^{2}}"


"\theta" = 30"\degree"


Equation of Force  "F = IBL sin \theta"


Force per unit length "\frac{F}{L} = IB sin \theta"


"\frac{F}{L} = 5\times 17\times sin 30\degree"


"\frac{F}{L} = 42.5 \frac{N}{m}"